OpenStudy (anonymous):

How many 1/3 are in 1/2?

OpenStudy (johnweldon1993):

So you would divide them... 1/2 --- = 1/2 x 3/1 = ? 1/3

OpenStudy (anonymous):

would it be 3/2?

OpenStudy (johnweldon1993):

Correct!

OpenStudy (anonymous):

awesome thank u

OpenStudy (johnweldon1993):

Anytime :)